MICROWAVE SPECTRUM OF PENTAFLUOROBENZONITRILE

\underline{S. D.SHARMA}, Spectroscopy Division, Bhabha Atomic Research Centre, Mumbai-400085, INDIA.

The microwave spectrum of pentafluorobenzonitrile has been reinvestigated in the light of subsequent studies of many fluorosubstituted benzenes. The rotational and centrifugal distortion constants obtained (in MHz) are Ã = 1029.027(24),B̃ = 759.956(14),C̃ = 437.554(14), Δ_{J} = −0/000096(13),Δ_{JK} = 0.00070(5),Δ_{K} = −0.00005(11),δ_{J} = −0.000016(7) and δ_{K} = −0.00044(3). The structural distortion of the benzene ring is on the expected lines with other fluorosubstituted benzene molecules.